OpenStudy (biohazard9064):

Please help Im stuck on the last one. Solve for x: 2 over 5 (x - 4) = 2x.

OpenStudy (anonymous):

\[\frac{2}{5}(x-4) = 2x\]

OpenStudy (anonymous):

Step : 1 - Multiply by 5 both the sides first.. Step : 2 - Then divide by 2 both the sides.. Step : 3 - Then, subtract \(x\) from both the sides. Step : 4 - Then, divide by 4 both the sides. Finally, you will get the value of \(x\)..

OpenStudy (biohazard9064):

so multiply 2 over 5 by 5

OpenStudy (biohazard9064):

and 2x?

OpenStudy (anonymous):

yes.. What you got after step 1?

OpenStudy (biohazard9064):

10 over 5

OpenStudy (anonymous):

Just do it, if you are wrong, I will correct, now try, don't hesitate..

OpenStudy (anonymous):

Can't you cancel 5 with 5 there?

OpenStudy (biohazard9064):

yes

OpenStudy (biohazard9064):

oh the 5 cancels that 5 leaving me with 2

OpenStudy (anonymous):

\[\frac{2}{5}(x-4) \times 5 = 2x \times 5 \\ \frac{2}{\cancel{5}}(x-4) \times \cancel{5} = 10x \\ 2(x-4) = 10x\]

OpenStudy (anonymous):

This is what you get after step 1.. Now what you get after step 2??

OpenStudy (biohazard9064):

then I divide by 2

OpenStudy (anonymous):

yeah do it..

OpenStudy (biohazard9064):

(x-4)=5x

OpenStudy (anonymous):

Wow, superb..!! :) Now (x-4) you can write it as : x - 4, without brackets too, no harm...

OpenStudy (anonymous):

\(x - 4 = 5x\)

OpenStudy (anonymous):

Now, subtract \(x\) from both sides which is our 3rd step. Do it..

OpenStudy (biohazard9064):

-4=5

OpenStudy (anonymous):

LHS is -4.. that is only good, what about RHS?

OpenStudy (anonymous):

\[x - 4 \color{red}{- x} = 5x \color{red}{-x} \\ \cancel{x} - 4 \color{red}{\cancel{-x}} = 5x \color{red}{- x} \\ -4 = 5x \color{red}{- x}\]

OpenStudy (anonymous):

Now tell me what is \(5x-x\) = ??

OpenStudy (biohazard9064):

5

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

I tell you by one more method.. Okay, take \(x =2\), now \(5x = 5 \times 2 = 10\)

OpenStudy (anonymous):

Is this good? I have taken the value of x = 2, x is any number there, it is variable, it can take any value, so I took its value as 2 just as an example...

OpenStudy (anonymous):

\(5x\) means \(5 \times x\) that is 5 is multiplied with \(x\)..

OpenStudy (biohazard9064):

sorry my computer froze and reset

OpenStudy (biohazard9064):

so it pasted my last message

OpenStudy (anonymous):

If I take x = 2, then \(5x\) becomes \(5 \times x = 5 \times 2 = 10\), this you got?

OpenStudy (biohazard9064):

yes

OpenStudy (anonymous):

So, you have :